Circular No. 6215

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

STELLAR OCCULTATIONS BY NEPTUNE I (TRITON) J. Elliot, C. Olkin, and H. Hammel, Massachusetts Institute of Tech-

nology (MIT), write: “Several observing teams successfully observed occultations of the star Tr148 (McDonald and Elliot 1995, A.J. 109, 1352) by Triton, including a central flash observed with the Infrared Telescope Facility (IRTF) at Mauna Kea simultaneously in visible wavelengths and in band K. Refined predictions just prior to the event were generated at MIT by J. Foust and S. McDonald using CCD observations made at Lick Observatory by J. McDonald, E. Dunham, and R. Stone; at Wallace Observatory by M. Mattei and Foust; and at the U.S. Naval Observatory (USNO) by A. Bosh, C. Dahn, and Stone. The occultation observations revealed that Tr148A has a fainter companion Tr148B: Δm (unfiltered CCD visible light) ∼ 1.5, separation ∼ 0′′.4. Approximate halflight times (i.e., the times of immersion and emersion, when the stellar flux was halfway between its unocculted and fully-occulted values) for the observed chords are given as minutes and seconds past 1995 Aug. 14^{d}07^{h} UT (instruments used unfiltered CCDs in visible light except where denoted K band): IRTF 3-m reflector, chord B, 33^{m}16^{s}–35^{m}10^{s} (Olkin, Hammel, and A. Cooray); Lick Crossley 0.9-m reflector, chord A, 36^{m}42^{s}–37^{m}57^{s} and chord B, 31^{m}08^{s}– 32^{m}44^{s} (J. McDonald and L. Young); Kuiper Airborne Observatory 0.9-m reflector, chord A, 38^{m}35^{s}–40^{m}14^{s} (Elliot and Dunham; WGS84 beginning and end positions: λ = 157^{o}30′35′′.64 west, ϕ = +32^{o}36′14′′.28, h = 13.00 km; 157^{o}42′26′′.28 west, +32^{o}32′36′′.48, 12.95 km); Lowell 1.8-m reflector, chord B, 30^{m}28^{s}–32^{m}22^{s} (M. Buie, L. Wasserman, and R. Millis); Wyoming Infrared Observatory 2.3-m reflector, chord A, 36^{m}13^{s}–37^{m}03^{s} (R. Hill, H. Reitsema, and R. Howell). Observations were also obtained with the Yerkes 1.0-m telescope (+ unfiltered CCD) by W. Wild, B. Drish, and J. Briggs; and at the Lick 1.0-m telescope (band K) by D. Rank, D. Gilmore, and J. Holbrook, but times are not yet available. Coverage at the Steward 2.3-m reflector by W. Hubbard and R. Marcialis was clouded out. Other observers of the event are encouraged to contact Olkin (colkin@astron.mit.edu).”

COMET 73P/SCHWASSMANN-WACHMANN 3 Total magnitude estimate: Aug. 19.46 UT, 12.9 (K. Kinoshita, Hi-

roshima, Japan, 0.28-m reflector + CCD; coma diameter 1′.1).

1995 August 28 Daniel W. E. Green